OpenStudy (legomyego180):

Need help with some algebra: 4x-2x^2=-12

OpenStudy (legomyego180):

Ive rewritten this as: \[x^2-2x-6=0\]

OpenStudy (photon336):

step one let's get all the terms to the same time. \[4x-2x^{2}+12 =0\] \[2(2x-x^{2}+6) = 0 \] \[2*(1/2)(2x-x^{2}+6) = 0*(1/2) \] \[2x-x^{2}+6 = 0 \]

OpenStudy (legomyego180):

cool ok

OpenStudy (photon336):

Yeah, do you know what to do next?

OpenStudy (legomyego180):

I know I could use the quadratic formula, but I dont think Im supposed to

OpenStudy (legomyego180):

My answer is in the form of \[1+\sqrt{7}\]

OpenStudy (photon336):

let me show you another way

OpenStudy (legomyego180):

It embarassing lol, this is for a calc II class and I forgot how to simplify

OpenStudy (photon336):

yeah so let's go off from where we started. \[2x-x^{2}+6 = 0\]

OpenStudy (legomyego180):

k

OpenStudy (photon336):

first thing i'm going to do is multiply everything by -1 because I like to have the first term -x^2 to be positive

OpenStudy (photon336):

\[-1(2x-x^{2}+6) = 0*(-1) => -2x+x^{2}-6\] \[x^{2}-2x-6 = 0\] now let's do something called completing the square. first thing we must do is get it in the form like this. \[x^{2}-2x = 6\] \[ax^{2}+bx = c\] now let's take (b/2)^2 and add to both sides \[x^{2}-2x+(\frac{ -2 }{ 2 })^{2} = 6+(\frac{ -2 }{ 2})^{2}\] \[x^{2}-2x+1 = 7\] now the next part is what we do here is interesting called completing the square we re-re-write this as the following. \[(x-1)^{2} = 7\] \[(x-1) = \pm~\sqrt{7}\] \[Thus~here~are~our~roots~(1~\pm~\sqrt{7})\]

OpenStudy (legomyego180):

Ah gotcha, I always hated completing the square for some reason. Something Im going to need to practice to remember. Thank you for the help!

OpenStudy (photon336):

let's check this answer. \[2(1+\sqrt{7})-(1+\sqrt{7})^{2}+6\] \[2x-x^{2}+6 => (2+ 2\sqrt{7})-(1-2\sqrt{7}-7)+6~ -8+8+2\sqrt{7}-2\sqrt{7} = 0\]

OpenStudy (legomyego180):

ok gotcha, so the you have to check the negative too?

OpenStudy (photon336):

both would give you the same answer .

OpenStudy (photon336):

lets me see about your other question

OpenStudy (legomyego180):

same question

OpenStudy (legomyego180):

Just thought it would help if you could see where I was coming from

OpenStudy (agent0smith):

So we can add basic algebra alongside fractions, in the list of things this calculus student doesn't know :P

OpenStudy (agent0smith):

@legomyego180 why do you not think you're supposed to use the quadratic formula? There's a reason it exists - it is essentially the same thing as completing the square, but takes way less time.

OpenStudy (photon336):

yeah exactly like you don't really have to do all that work I did

OpenStudy (agent0smith):

At some point, someone realized "hey, completing the square every time like this is tedious, quite tedious indeed, why don't we find a way to make this much less a suckfest"

OpenStudy (legomyego180):

I didnt think it would give you the same answer

OpenStudy (legomyego180):

Yea its a long and complicated story about how I took precal and algebra 4 years ago and now im taking calculus and forgot everything blah blah blah

OpenStudy (agent0smith):

Haven't you ever seen the derivation of the quadratic formula? It literally results from completing the square on the general form of a quadratic, ax^2 + bx + c = 0.

OpenStudy (legomyego180):

I dont think so

OpenStudy (agent0smith):

And I don't blame you for forgetting how to complete the square. It does kinda suck. Forgetting how to deal with fractions, though... :P